# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

### v3

@@ 7 Conclusion
This work introduces the concept of antifragility by comparing it with two other terms, i.e., robustness and resilience, which are commonly used in transportation and traffic control strategies. Through reviewing previous research on applying RL algorithms in traffic control to achieve robust or resilient designs, we explore whether and how these properties can be induced with RL. Building on this concept, we aim to induce antifragility in perimeter control by modifying the state definition and reward function based on a state-of-the-art RL-based perimeter control algorithm. We incorporate the first and second derivatives — representing the rate of change and curvature of traffic states — to provide the RL algorithm with richer information. To mitigate the potential oscillations caused by integrating derivatives, a damping term is introduced to stabilize the computed actions. Furthermore, a redundant overcompensation term in the reward function strengthens the system’s antifragility against disruptions.
@@
We conducted comprehensive experiments to compare our proposed antifragile RL-based perimeter control approach against three other methods: NC, MPC, and the state-of-the-art RL-based method as the baseline. This evaluation was performed on two levels—disruptions with static and incremental magnitudes to validate proto-antifragility and progressive antifragility, respectively. In addition, we examined two distinct manifestations of disruptions: demand disruptions and supply disruptions. Uncertainties in the magnitude of disruptions are also considered. The results demonstrated the effectiveness of our proposed method based on superior performance regarding lower TTS. More importantly. The proposed algorithm showcased different levels of antifragility under static and incremental disruptions. We put forward a novel method for quantifying antifragility by computing distribution skewness. The proposed RL-based antifragile algorithm exhibits the lowest skewness among all the methods examined, indicating its progressive antifragility property compared to the other examined methods.
@@
In conclusion, this study is the first of its kind to pioneer the application of antifragility in the operation of transportation networks, improving performance under unforeseen disruptions with learning-based algorithms. In addition, it introduces a new possibility for evaluating system operation under disruptive conditions. Moreover, the concept is generic to be extended to other traffic control systems and potentially to systems that are subject to disruptions in general.
@@ 10 Declaration of Competing Interest
Appendix A Graphical illustration of robustness, resilience, and antifragility
Fig. 13 graphically illustrates the differences in robustness, resilience, proto-antifragility, and progressive antifragility. In brief, robustness is about the system’s resistance to minor disturbances, whereas resilience represents the ability to recover from major disruptions. Proto-antifragility enables the system to adapt to disruptions with magnitudes falling within a similar range, while the performance of a progressive antifragile system exhibits a concave response with linearly growing disruptions, either due to the inherent properties of the system or through antifragile control strategies.

### v4

@@ 7 Conclusion
As disruptions are ubiquitous in urban transportation systems and are expected to increase continuously with population growth and ongoing urbanization, this work introduces the novel concept of antifragility into traffic control to tackle the growing trend of disruptions and the fragile nature of road transportation networks. First, from a terminology perspective, antifragility is compared with other related concepts commonly applied in transportation, including robustness, resilience, reliability, and adaptiveness. Through reviewing previous research on applying RL algorithms in traffic control to achieve robust or resilient traffic control systems, we explore whether and how these properties can be induced with RL. Building on such concepts, we aim to induce antifragility in perimeter control by modifying the state definition and reward function based on a state-of-the-art RL-based perimeter control algorithm. We incorporate the first and second derivatives, representing the change rate and the curvature of traffic states, to provide the RL algorithm with richer and reliable information under disruptive events. To mitigate the potential oscillations caused by integrating derivatives, a damping term is introduced to stabilize the computed actions. Furthermore, a redundant overcompensation term in the reward function strengthens the system’s antifragility against disruptions. In addition to the performance gain and antifragility, another contribution of this work is the consideration of real-world observability constraints, complementing the drawback of the current state-of-the-art algorithms of utilizing hardly observable measurements.
@@
We conducted comprehensive experiments to compare our proposed antifragile RL-based perimeter control approach against three other methods: MPC, MPC-MHE, and an RL-based algorithm. Two distinct manifestations of disruptions are examined, i.e., demand and supply disruptions. Uncertainties in the disruption magnitude are also considered. The results first demonstrate the effectiveness of our proposed antifragile RL-based algorithm, through its superior performance in terms of lower TTS. More importantly, by putting forward a novel method for quantifying antifragility through computing its performance distribution skewness, we confirm that the proposed algorithm exhibits greater antifragility with the lowest skewness among all the methods examined, delivering increasingly better performance as the magnitude of disruptions grows. Finally, the performance and antifragile properties are studied with limited real-world observability, showing that a certain degree of performance can be traded for using only accessible sensor measurements in reality.
Several limitations of the current study need to be addressed as well. For instance, measurements and model uncertainties are commonly studied when designing perimeter control algorithms to validate their robustness. However, to uphold the clarity of this work, that is, studying antifragility instead of robustness, and also due to our introduction of uncertainties across the episodes with incremental disruptions, we disregard the uncertainties within each episode, but we acknowledge the common presence of uncertainties in the real urban networks. Furthermore, as this work primarily works on numerical simulation, the real-world operation together with the proposed antifragile RL-based algorithm can be better validated with microsimulation software, such as SUMO, which is subject to our future work.
@@
In conclusion, this study is the first of its kind to pioneer the application of antifragility in the operation of transportation systems, continuously improving the performance of a network under unforeseen disruptions with learning-based algorithms. Additionally, it introduces a new paradigm for evaluating system operation under disruptive conditions. Moreover, the concept is generic to be extended to other traffic control systems and potentially to any system subject to growing disruptions.
@@ 10 Declaration of Competing Interest
Appendix A Graphical illustration of concepts related to antifragility
Fig. 16 graphically illustrates the differences in robustness, resilience, adaptiveness, and progressive antifragility. In brief, robustness is about the system’s resistance to minor disturbances, whereas resilience represents the ability to recover from major disruptions. Adaptiveness enables the system to adapt to disruptions with magnitudes falling within a similar range, while the performance of a progressive antifragile system exhibits a concave response with linearly growing disruptions, either due to the inherent properties of the system or through antifragile control strategies.